Acme Oscillators is considering an investment project that has the following rather unusual cash flow pattern. Year Cash Flow 0 ​$100 1 −460 2 791 3 −602.6 4 171.6 a. Calculate the​ project's NPV at each of the following discount​ rates: 30%​, 40​%, 50​%. b. What do the calculations tell you about this​ project's IRR? The IRR rule tells managers to invest if a​ project's IRR is greater than the cost of capital. If Acme​ Oscillators' cost of capital is 8​%, should the company accept or reject this​ investment? c. Notice that this​ project's greatest NPVs come at very high discount rates. Can you provide an intuitive explanation for that​pattern?
Answer #1

1.
NPV at 30%=100-460/1.3+791/1.3^2-602.6/1.3^3+171.6/1.3^4=0

2.
NPV at 40%=100-460/1.4+791/1.4^2-602.6/1.4^3+171.6/1.4^4=0.0624739691794716

3.
NPV at 50%=100-460/1.5+791/1.5^2-602.6/1.5^3+171.6/1.5^4=0.237037037036991

4.
Approximate IRR is 30%
Accept

Actual IRR is 0%, 10%, 20%, 30%

5.
As outflows are happening at later time periods, higher discount rate lowers the present value of cash outflows thus improving the NPV